OpenStudy (anonymous):

Determine the critical numbers of the given function and classify each critical point as a relative maximum, relative min or neither. f(t)= t/(t^2+3) I can get to t/(t^2+3) I need to solve for 0 to find the intervals of increase and decrease. I have forgotten how to solve for 0 with this problem.

OpenStudy (anonymous):

You got nowhere.

OpenStudy (anonymous):

I know i need the 1st derivitive. Which I got t(2t) - (t^2+3)(1) / (t^2+3)^2 Is that much correct?

OpenStudy (amistre64):

f'(x) = [ BT' - B'T ] / B^2

OpenStudy (amistre64):

(t^2 +3)(1) - (t)(2t) ---------------- (t^2 +3)^2

OpenStudy (amistre64):

any second opinions? :)

OpenStudy (nowhereman):

I agree, but _please_ use the equation-editor: \[f'(t) = \frac{t^2 + 3 - 2t\cdot t}{(t^2+3)^2}\]

OpenStudy (amistre64):

t^2 +3 -2t^2 = 0 -t^2 +3 = 0 -t^2 = -3 t^2 = 3 t=+-sqrt(3)

OpenStudy (anonymous):

I have that. I just have the top reversed.

OpenStudy (nowhereman):

and you were missing parentheses

OpenStudy (amistre64):

I cant get that equation editor to work right .....

OpenStudy (nowhereman):

you can simply enclose the latex-code in \ [ and \ ].

OpenStudy (amistre64):

latex is for painting houses ;)

OpenStudy (anonymous):

amistre64 is right: \[f'(x)= \frac{3-t^2}{(t^2+3)^2}\]

OpenStudy (amistre64):

something likethis? testing \[45\Omega -\sin(45) -\infty\]

OpenStudy (amistre64):

I do better after a nap :)

OpenStudy (nowhereman):

sleep well :-)